OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal...

OpenStudy (abbycross167):

Mary can row at the rate of 4 miles per hour in still water. Rowing downstream in a river, she can travel 3 times as far in 1 hour as she can rowing upstream. What is the rate of flow in the river?

OpenStudy (mathmate):

Let x=speed of current Mary's downstream speed = 4+x Mary's upstream speed = 4-x We know that 4+x is 3 time 4-x, or in mathematics equation: \(\Large 4+x=3(4-x)\) Solve for x, taking care to distribute multiplication when removing brackets on the right hand side.

OpenStudy (abbycross167):

so x =2

OpenStudy (abbycross167):

@mathmate

OpenStudy (mathmate):

X=2 is correct, well done! :)

OpenStudy (abbycross167):

and thats it? thats all I have to do for this question?

OpenStudy (abbycross167):

thank you so much for the help!! :)

OpenStudy (mathmate):

Yep, math is not as hard as many say! :)